oncology nursing - calculate the ANC - WBC 700/mm3, polys- 11%, bands-8 a) 567/mm3, b) 133/ mm3, c) 13/ mm3. d) 77/ mm3
Answer #1

Ans: b) 133 /mm3

Calculation:

Given values: WBC 700/mm3, polys- 11%, bands-8

(assuming the bands' value given in the question is in %)

ANC = WBC * (polys% + bands%) / 100

ANC = 700 * (11+8)/100

ANC = 133/mm3